OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x,y). 2x+3y=9 y=2x+13

jhonyy9 (jhonyy9):

so there are 2x+3y=9 y=2x+13 using substitution method you need just substitute the y value from second equation in place of y inside first equation ihan solve this and will get the value of y so this will need to replace it inside second equation and you can calcule the value of x so than will get the right answer (x,y)

jhonyy9 (jhonyy9):

little help from second equation y=2x+13 so this substitute in place of y inside first equation and will get 2x+3(2x+13)=9 now just you need to solve this for x 2x+6x+39=9 8x=9-39 8x=-30 x=-30/8 x=-15/4 y=2x+13 y=2(-15/4) +13 y=-15/2 +13 y=(-15+26)/2 y=11/2 (-15/4 ,11/2)